OpenStudy (anonymous):

find solutions in interval [0,2pie) for csc^2x=cotx+1

OpenStudy (anonymous):

x=pi/8

OpenStudy (anonymous):

thats not a choice. do you want me to list the choices ?

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

sorry pi/4

OpenStudy (anonymous):

or 45 degrees

OpenStudy (anonymous):

I mean solution is cosx=sinx

OpenStudy (anonymous):

A. pie/4 , 3pie/4 , 4pie/3 B. pie/2 , 3pie/2 C. 0 , pie/2 , pie , 3pie/2 D. pie/4 , pie/2 , 5pie/4 , 3pie/2 E. pie/4 , 4pie/3 , 11pie/6

OpenStudy (anonymous):

use the identity\[\csc^2 x = 1 + \cot^2 x\]\[\csc^2 x = \cot x + 1\]\[1 + \cot^2 x = \cot x + 1\]\[\cot^2 x - \cot x = 0\]\[\left( \cot x \right)\left( \cot x - 1 \right) = 0\]can you continue after this ?

OpenStudy (anonymous):

not really. i don't understand how that gets any of the answers

OpenStudy (anonymous):

\[\cot x = 0\]or \[\cot x - 1 = 0\]\[\cot x = 1\]find all values of x within the interval [0,2π) that satisfies cot x = 0 or cot x = 1

OpenStudy (anonymous):

OHH thanks !!